OpenStudy (anonymous):

Simplify b^3 × b^-3 for b = 5.

OpenStudy (anonymous):

b^3 is the same as b * b *b b^-3 is the same as 1/b^3 = 1/b * 1/b * 1/b so what you have is: \[(b * b * b) * (\frac{ 1 }{ b } * \frac{ 1 }{ b } * \frac{ 1 }{ b} )\] which is the same as \[\frac{ b * b * b }{ b * b*b} \] all of the b's cancel out to become 1 regardless of what value b is: Example: \[\frac{ 5 * 5 * 5 }{ 5 * 5 * 5 } = \frac{ 125 }{ 125 } = 1\]

OpenStudy (anonymous):

Thank you :)